Determine the​ upper-tail critical value of F in each of the following​ one-tail tests for a claim that the variance of sample 1 is greater than the variance of sample 2.

A) a = 0.05, n1= 10,n2= 13 B) a= 0.025, n1 = 10, n2 = 13
Answer #1

a)

df1 = n1 - 1 = 10- 1 = 9

df2 = n2 - 1 = 13 - 1 = 12

From F Critical value table,

F critical value at 0.05 significance level with 9 and 12 df = 2.796

b)

df1 = n1 - 1 = 10- 1 = 9

df2 = n2 - 1 = 13 - 1 = 12

From F Critical value table,

F critical value at 0.025 significance level with 9 and 12 df = 3.436